### v1

# These authors contributed equally to this work

Abstract
Analog bolometers based on temperature-dependent phase-transition materials such as vanadium oxide (VOx) and barium titanate (BTO) represent the state of the art in uncooled infrared detectors. Recently, the first room-temperature spintronic Poisson bolometer based on magnetic tunnel junctions (MTJs) was proposed and demonstrated as a promising infrared detector. Unlike conventional bolometers, the spintronic Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, where the response is governed by resistance fluctuations arising from thermally activated magnetization transitions. Spontaneous transitions between two metastable magnetic states occur even in the absence of incident light, and the transition probability increases under illumination. In this work, we experimentally study the statistical properties of the optical response of the spintronic Poisson bolometer under illumination. We demonstrate that transitions in spintronic Poisson bolometers, both in the absence and presence of light, exhibit Poissonian behavior, with transition rates and interarrival times modulated by incident radiation. Under illumination, we observe a 153% increase in the count rate accompanied by a 70% reduction in interarrival time. These results establish spintronic Poisson bolometers as a promising platform for probabilistic, high-speed, and high-sensitivity infrared detection at room temperature.
preprint: AIP/123-QED
I Introduction
Bolometers are thermal detectors in which incident light causes a temperature-induced resistance change that is read out electrically YADAV2022113611. Infrared bolometers are essential for emerging applications in thermal imaging, autonomous navigation, astronomy, and time-of-flight sensing Bao2023; Lahiri2012; Sivaprakasam2020; Kaushal2017; anderson2023advancements; rodriguez2023remote; PhysRevX.14.041005. State-of-the-art uncooled bolometers are analog detectors based on temperature-dependent phase transition of materials such as vanadium oxide (VOx) and barium titanate (BTO) yeh2020performance; scott2022sensing. Other uncooled bolometers are based on semiconductor and metallic materials with temperature-dependent resistance, including amorphous silicon and platinumdoi:10.1021/acsphotonics.9b01198; doi:10.1021/acs.nanolett.1c02972; doi:10.1021/acs.nanolett.3c03076. In VOx bolometers, major noise sources are Johnson noise, thermal noise, generation–recombination and $1/f$ noise philipose2023analyzing; 10.1063/5.0049633; du2013graphene. Similarly, the dominant noise in barium titanate (BTO) bolometers and graphene bolometers is Johnson noise Hanel:61; du2013graphene; PhysRevB.96.165431. All of the uncooled bolometers aforementioned are analog detectors that produce a continuous voltage or current proportional to the incoming signal, and are dominated by Gaussian noise, as shown in Fig. 1(a).
Recently, the first uncooled spintronic Poisson bolometer, called the spintronic ultrafast nanoscale bolometer (SUN bolometer) leif_nanolett was proposed. Unlike analog bolometers, the spintronic Poisson bolometers convert incoming signals into discrete counting events and are dominated by Poissonian noise, as shown in Fig. 1(b). In spintronic Poisson bolometers, magnetic tunnel junctions (MTJs) are engineered with an energy barrier of the order of thermal infrared energy at room temperature. It exploits switching between two discrete magnetization states by intrinsic thermal fluctuations and nanoscale heating induced by incident radiation. These digital probabilistic transitions in the absence of light are referred to as dark counts, and when under illumination, are bright counts. It was demonstrated that the count rate increases due to the incident light, validating the device’s response through the Néel-Arrhenius law PhysRevB.104.094433. The spintronic Poisson bolometer achieved a noise equivalent temperature difference (NETD) of 103 mK at a 25 Hz frame rate, and high readout speeds of 10 MHz, which is faster than commercial vanadium dioxide detectors. The high sensitivity and the high readout speed make it a promising candidate for state-of-the-art infrared detection.
In other studies in the field, stochastic MTJs have been utilized for random number generation, probabilistic computing, and neuromorphic computing choi2014magnetic; joshi2020mtj; zhang2021recent; maciel2019magnetic; zink2022review. The stochastic magnetization transitions, known as magnetic random telegraph noise (RTN), exhibit relaxation times that follow an exponential distribution, consistent with Poisson event statistics hayakawa2021nanosecond; 2021relaxation. Similarly, detection-based studies have shown exponentially distributed dark-count interarrival times leif_nanolett. Notably, previous studies have examined the statistics of MTJ transitions only in the absence of incident light.
In this work, we extend the study to the optical response of the spintronic Poisson bolometer under incident radiation. The spintronic Poisson bolometer operates in a noise-dominated probabilistic regime, where the response is governed by random resistance fluctuations arising from thermally activated magnetization switching. We demonstrate that both dark-count and bright-count statistics in spintronic Poisson bolometers are well described by the Poisson distribution, with the mean count rate modulated by incident radiation. Changes in the mean number of counts indicate enhanced thermal switching between the two magnetization states in the MTJ sensing layer, serving as a key to quantifying incident radiation. The spintronic Poisson bolometer thus offers a fundamentally different detection paradigm based on an indirect, thermally mediated light–matter interaction, leading to Poisson statistics that are independent of the light source statistics.
The Poissonian nature of these counts enables time-resolved measurements with high fidelity and superior signal-to-noise ratio (SNR) compared to analog systems under identical conditions. The spintronic Poisson bolometer also exhibits distinct statistics compared to other digital detectors due to the absence of after-pulsing WANG2019202 and its intrinsic Poisson response, irrespective of light statistics. We also investigate the device’s non-probabilistic response. Excitation with a pulsed laser produces current transients, providing insights into heat transfer and transport within the device. This systematic experimental investigation of the response statistics of spintronic Poisson bolometers under illumination advances our understanding of the underlying device physics and guides optimization of sensitivity, response time, dynamic range, and spectral response. Overall, we establish the spintronic Poisson bolometer as a promising platform for probabilistic, high-speed, and high-sensitivity infrared detection at room temperature.
Figure 1: (a) In analog detectors, thermal noise and Johnson noise follow a Gaussian distribution as they stem from thermal agitation of a large number of electrons. (b) In digital detectors, dark counts follow a Poisson distribution as they stem from discrete independent events happening at a constant mean rate, such as thermal excitations and quantum tunneling effects in single photon detectors (SPDs), or magnetization flips governed by exponentially distributed interarrival times in spintronic Poisson bolometers (SPB). (c) Device structure. The main structure has five layers: a transduction layer, a magnetic free layer, a tunnel barrier, a magnetic fixed layer, and a synthetic antiferromagnetic layer (SAF). The magnetization M of the readout layer is pinned by the SAF. The magnetization of the sensing layer flips between M1 and M2 probabilistically due to heat. (d) When no light is incident, natural heat in the device causes discrete transitions between these two magnetization directions, which we refer to as dark counts. When light is incident, a hot spot is formed in the transduction layer, which thermalizes through a magnetic tunnel junction (MTJ). This increases the probability of transition in the MTJs sensing layer, leading to an increased count rate, which is read out through the device’s resistance. (e) Left: Optical image of the device. Right: SEM image of the MTJ nanopillar before Au deposition.
II Results and Discussion
The spintronic Poisson bolometer studied here consists of five layers: a transduction layer at the top, a magnetic sensing (free) layer, a nonmagnetic insulating tunnel barrier, a magnetic readout (fixed) layer, and a synthetic antiferromagnetic (SAF) layer, as shown in Fig. 1(c). The latter four layers form a magnetic tunnel junction (MTJ) that exhibits tunnel magnetoresistance. The magnetization of the readout layer is pinned by the underlying SAF layer, which stabilizes an out-of-plane (perpendicular) magnetization in the fixed layer. The sensing layer possesses two stable magnetization states (M1 and M2). Due to the presence of spin-polarized current in the device, the resistance depends on the sensing layer’s magnetic orientation.
Even without illumination, thermal fluctuations induce spontaneous transitions between the two sensing-layer states, producing dark counts. Illumination generates a hot spot in the transduction layer, which thermalizes through the MTJ stack and increases the probability of transitions in the sensing layer. This results in an increased count rate observed through MTJ resistance readout (Fig. 1(d)). The behavior of these transitions follows the Néel–Arrhenius law for thermally activated switching2021relaxation:
$\tau_{M_{1},(M_{2})}=\tau_{0}\exp\left(\frac{E_{b}}{k_{B}T}\right)$ (1)
where $\tau_{0}$ is the attempt time, $k_{B}$ is Boltzmann’s constant, and $T$ is the temperature, and $E_{b}$ is the energy barrier:
$E_{b}=\frac{MV(H_{k}\mp H)}{2}$ (2)
where $M$ is the magnetization, $H_{k}$ is the magnetic anisotropy field, V is the volume of the sensing layer, and H is the magnetic field present in the sensing layer. Fig. 1(e) shows an optical image of the fabricated spintronic Poisson bolometer device and an SEM image of the MTJ nanopillar before Au deposition.
We next present a general model for detector signal and response time. For analog bolometers such as VOx, the detector signal is an output voltage, whereas for the spintronic Poisson bolometer, the signal is the count rate. The relationship between incident power and detector signal is quantified through the temperature increase of the bolometer, which depends on the thermal properties:
$\Delta T=\frac{\varepsilon P/A}{\left(G_{\mathrm{th}}^{2}+\omega^{2}C_{\mathrm{th}}^{2}\right)^{1/2}}$ (3)
where $\Delta T$ is the temperature change of the detector, $\varepsilon$ is emissivity, $P$ is the incident radiant power, and $A$ is the effective detector area, $G_{\mathrm{th}}$ is the thermal conductance, $C_{\mathrm{th}}$ is the thermal capacitance, and $\omega$ is the angular modulation frequency 8845784.
The temperature-to-signal conversion is characterized by-
$k=\frac{dV}{dT}$ (4)
which reflects the sensitivity of the detector output to temperature change.8845784 For VOx, $k$ is constant over the operating range, whereas for the spintronic Poisson bolometer, the change in count rate follows the Neel-Arrhenius relation leif_nanolett. For small changes in temperature, the count rate increase is
$dN=\frac{2}{\tau_{0}}\exp(\frac{-E_{b}}{k_{B}T})*\frac{E_{b}}{k_{B}T^{2}}*dT$ (5)
where $dN$ is the increase in counts due to a temperature rise $dT$, $\tau_{0}$ is the relaxation time, $E_{b}$ is the energy barrier, $k_{B}$ is the Boltzmann constant (Supplementary I). Thus, in both conventional analog bolometers and digital spintronic Poisson bolometers, the signal is linearly dependent on temperature.
The detector response time is given by
$\tau_{total}=\tau_{diffusion}+\tau_{thermal}+\tau_{intrinsic}$ (6)
where $\tau_{\text{diffusion}}$ is the diffusion time, $\tau_{\text{thermal}}$ is the thermal time constant, and $\tau_{\text{intrinsic}}$ represents intrinsic system dynamics as magnetic spin switching. For VOx bolometers, the response time is dominated by thermal time constant (10 ms), whereas for the spintronic Poisson bolometer, it is governed by the diffusion time (2.2-3.5 ns)Abdullah2023; leif_nanolett. Because the spintronic Poisson bolometer is a probabilistic Poisson detector, deterministic readout of a transition is further limited by approximately 4.6 times the relaxation time (0.1-10 $\mu s$) (Supplementary II). Overall, the spintronic Poisson bolometer is faster than $VO_{x}$ by three orders of magnitude, making it a strong candidate for high-speed detector applications.
To study the spintronic Poisson bolometer response to incident light over a range of pulse widths, powers, and wavelengths, we used three different laser sources: (1) a continuous-wave (CW) 405 nm laser, (2) a 808 nm CW laser with power modulation by an arbitrary waveform generator, and (3) a 405 nm pico-second laser with a 20 ps pulse width. A schematic of the corresponding power-time profiles is shown in Fig. 2 (a).
II.1 Device Response to 808 nm Modulated Laser
Figure 2: (a)Schematic of the power-time relationship of 3 laser sources. We study the device response to 3 types of laser sources. The first laser is a continuous-wave (CW) 405 nm laser. The second laser is a modulated 808 nm CW laser. The power output of this laser is modulated by an arbitrary waveform generator (AWG). The third laser is a 405 nm pico-second laser that has a pulse width as small as 20 ps. (b) Raw readout signal. The orange curve is the laser trigger signal. (c) Time correlation. We divide each period into 100 bins (100 ns per bin) and calculate the probability of a count happening in each bin. The red dashed line marks the time of the trigger of one of the periods when it turns the laser on, which occurs at t = 3233.6 ns. The system has around 80 ns latency between the laser trigger and light incidence on the device. We can see a stabilized count probability increase at t = 3366.7 ns, which is around 53 ns after the light is incident when considering the latency. (d) Count statistics. We divide the data into 20 $\mu s$ time bins and calculate the number of counts happening in each time bin. For both laser-on and off regions, the number of counts follows a Poisson distribution. With the fitted mean number of counts $\lambda$, we can calculate the mean count rates for laser off and on regions, which are 3.373285 Mcps and 3.95123 Mcps, respectively. This shows around a 17% increase in the count rate due to incident light.
[…]
Figure  2 (b) shows the device response to the 808 nm modulated laser. Counts are superimposed on the non-probabilistic voltage signal. When the laser is ON, we observe an increase in device resistance due to laser heating of the electrical contacts as well as an increase in count rate. To understand the statistics of ON and OFF regions, we divide each period into 100 bins (100 ns per bin) and the probability of detecting a count in each bin is computed over 6000 periods, as shown in Fig. 2 (c). The dotted line indicates the time light was incident on the device, and the red line is the voltage of the device averaged over 6000 periods. The probability of a count remains constant over both ON/OFF regions, indicating stochastic behavior which is characteristic of Poisson statistics. When the laser is turned ON, we observe that the probability of a count increases 53 ns after the light is incident (Supplementary).
We then separate the data into laser-ON and laser-OFF regions and study the probability distribution of the number of counts $P(N)$ for each region. We divide the data into 100 time bins and calculate the number of counts that occur in each time bin as shown in Fig. 2 (d). We observe that for both laser OFF and laser ON cases, the count distribution follows a Poisson distribution van1992stochastic:
$P(N)=\frac{\lambda^{N}}{N!}e^{-\lambda},N=0,1,2,$$\infty.$ (7)
From the Poisson fit of the probability distribution in Fig. 2 (d), the mean count rates for laser off and on regions are calculated to be 3.3 Mcps and 3.9 Mcps, respectively. This shows around a 17% increase in the count rate due to incident light. The shifted mean number of counts manifests increased temperature and energy in the sensing layer, and thus increased probability of transitions according to equation 5. These results suggest that the spintronic Poisson bolometer demonstrates Poissonian statistics for both dark counts and bright counts with incident radiation.
Figure 3: (a) Probability histogram of detector counts over $3.8~\mu\text{s}$ intervals under 405 nm CW laser illumination. The count statistics follow a Poisson distribution, as shown by the black fit curve. When the laser is off, the mean count is $\lambda$= 17; under illumination, the mean increases to $\lambda$ = 43, indicating light-induced switching events. (b) Voltage state distributions under dark and illuminated conditions show two distinct peaks corresponding to the parallel (low-resistance) and antiparallel (high-resistance) magnetization states. (c) Histogram of relaxation times spent in the parallel (low-voltage) state before transitioning to the antiparallel (high-voltage) state. The distribution follows an exponential decay, consistent with Poissonian statistics. Illumination reduces the mean relaxation time by 66%, increasing the switching rate. Inset: Bright count rate versus laser intensity shows a significant increase in counts in response to higher optical power. (d) Histogram of relaxation times in the antiparallel state before switching to the parallel state. Like the parallel case, the distribution follows an exponential profile, with mean relaxation time decreasing by 70% under illumination. Inset: Field dependence of both dark and bright count rates shows a maximum near H=1 mT, corresponding to the field strength where the energy barrier between magnetic states is minimized and switching is most probable.
II.2 Device Poissonian Response to 405 nm CW Laser
In this section, we examine the spintronic Poisson bolometer’s response to a 405 nm CW laser of intensity of $0.18\mathrm{mW}/\mu\mathrm{m}^{2}$ on the device and analyze the statistical behavior of both dark and bright counts, assessing the count statistics. The device registers a dark count rate of  4.5 Mcps. When illuminated with a 405 nm CW laser, the count rate increases to  11.4 Mcps—a 153 % increase. The gold top contact of the device absorbs approximately $60\%$ of incident light at 405 nm, compared to $1\%$ absorption at 808 nm. Thus, there is a higher count rate increase under a 405nm laser illumination than under 808nm, indicating a strong wavelength dependence of the statistical response.
Figure 3(a) presents histograms of detector counts over $3.8~\mu\text{s}$ intervals under dark and illuminated conditions. Both distributions follow Poissonian fits, with mean number of counts shifting from $\lambda$ = 17 (dark) to $\lambda$ = 43 (bright). This demonstrates that incident photons increase the probability of stochastic switching in the magnetic tunnel junction (MTJ), consistent with thermally activated transitions. Figure 3(b) shows the voltage state probability distributions for both conditions. Each case exhibits two distinct peaks corresponding to the parallel (low resistance) and antiparallel (high resistance) magnetization states of the sensing layer. Under illumination, a voltage shift is observed as a result of changes in magnetoresistance, providing further confirmation of light-induced modulation.
Figure 4: (a) For a 405 nm picosecond laser, the laser-induced response is observed in the spintronic Poisson bolometer readout when light is incident on or near the device. (b) We observe different behavior with the 808 nm laser, where the photocurrent response does not reduce further away from the device. Inset is the reflection spectrum of the Au top contact. (c)The pulse width dependence is studied by varying the duty cycle of the 808 nm modulated laser. We observe that for larger pulse widths, the photocurrent is reduced, which signifies power dependence. (d) The peak power dependence of the laser-induced voltage is linear at low powers and saturates at high powers.
To further investigate the switching dynamics, we analyze the relaxation time, i.e., the time between consecutive magnetization transitions. Figures 3(c) and 3(d) present histograms of relaxation times for the parallel and antiparallel states under both laser-off and laser-on conditions. In both cases, the distributions exhibit exponential decay, consistent with a Poisson process, as described by:
$P(\tau)=\frac{1}{\mu}e^{-\frac{1}{\mu}\tau}$ (8)
where, $\mu$ is the mean relaxation time. Under illumination, $\mu$ decreases significantly: from $121\text{ns}$ to $41\text{ns}$ in the parallel state, and from $266\text{ ns}$ to $81\text{ns}$ in the antiparallel state, resulting in an increased switching rate. This confirms that optical excitation enhances the probability of transition events, and is consistent with Néel–Arrhenius’ law (see Eq. 1), where the switching rate increases as the energy barrier decreases via optical heating that increases the temperature of the sensing layer.
The inset of Fig. 3(c) demonstrates that as the laser intensity increased from 0 to $0.18\,\frac{\mathrm{mW}}{\mu\mathrm{m}^{2}}$, the count rate nearly triples, from 4.5 Mcps to 11.4 Mcps—approaching saturation at higher intensities. Meanwhile, the inset of Fig. 3(d) shows that both dark and bright count rates follow a Gaussian dependence on the applied magnetic field, peaking near H = 1 mT where we operate our device. At this field bias, the energy barrier is minimum, leading to an increased transition rate. These results confirm the field-tunable nature of Poissonian distributed thermal switching events in the spintronic Poisson bolometer under 405 nm illumination, with optical modulation of both count rate and relaxation dynamics.
II.3 Laser Induced Current in MTJ-based detector due to a pulsed picosecond laser
Finally, we investigate the response of the spintronic Poisson bolometer to nanosecond and picosecond pulsed lasers. When the laser pulse is incident on the device, the absorbed light produces phonons and hot electrons. The top contact of the device is made up of 75 nm Au, which is known to induce bulk photocurrent due to pulsed laser. Since the device has an electrical readout, this photocurrent response is read out in addition to the MTJ tunnel magneto-resistance. We studied the induced-current response due to a 405 nm picosecond laser with a pulse width of 20 ps (Figure 4a) and an 808 nm modulated laser with a nanosecond (200-1000 ns) pulse width (Figure 4b) by analyzing the average voltage of the response. With high power of incident radiation, we observe a photocurrent in the metallic Au top contact.
The positional dependence of the photocurrent is studied for the two laser signals. For the 405 nm ps laser, we observe that the current is highest when light is incident on top of the device, and it reduces significantly when light is incident away from the device. This implies the presence of hot electron effects. Additionally, it takes the same time to get to the peak current for different powers of light, while the relaxation time depends on the power of incident radiation. For 808nm, we observe that the current induced is largely position-independent, indicating the presence of hot electron effects. The effect of the laser on the top and bottom contacts was also studied, with no major differences found in voltage, indicating the absence of thermo-voltaic effects. The presence of hot-electron effects at 405 nm can be explained by differences in the reflection spectrum. For 405 nm, Au is significantly absorptive (reflection=40%) while at 808 nm Au is highly reflecting (reflection=99%).
The induced current is studied for different pulse widths (Figure 4c). For smaller pulse widths, since the peak power of the laser is higher, the induced current is higher. The induced-current duration is smaller than an 800 ns pulse width, as it is a non-equilibrium effect. The power dependence is also analyzed (Figure 4d). The induced current is found to be linearly proportional to low incident powers and has a saturation at higher powers. Thus, we demonstrate a non-probabilistic response with ultrafast pulsed light and show a response dominated by photocurrent and hot-electron effects.
III Conclusion
In conclusion, we have experimentally demonstrated that spintronic Poisson bolometers exhibit Poissonian response statistics under both dark and illuminated conditions, independent of the statistics of the incident light source. Time-resolved measurements confirm that both dark and bright counts follow Poisson distributions, with interarrival times exhibiting exponential behavior. This validates the indirect thermally activated switching mechanism governed by the Néel-Arrhenius model. These findings establish spintronic Poisson bolometers as a new class of digital probabilistic infrared detectors with high-speed, high-sensitivity, room-temperature operation. By leveraging their inherent Poisson statistics, spintronic Poisson bolometers present a promising platform for next-generation time-resolved infrared detection and effectively address the limitations of current state-of-the-art analog detectors.
Acknowledgements.
This work is partially supported by the Elmore Chaired Professorship of Purdue University.